# On the identifiability of causal graphs with multiple environments

### v1

@@ C.2 Structural causal model identifiability from observational data
where $B\in\mathbb{R}^{d\times d}$ collects the coefficients expressing each $X_{i}$ as a linear function of its parents plus a disturbance $S_{i}$. With mutually independent, non-Gaussian noise terms, the model is identifiable; this is known as the Linear Non-Gaussian Acyclic Model (LiNGAM) (Shimizu et al., 2006).
@@
An Additive Noise Model (ANM) (Hoyer et al., 2008; Peters et al., 2014) defines each causal variable as a function of (potentially) nonlinear mechanisms and an additive noise contribution:
@@
The most general class with known sufficient conditions for identifiability of the graph is the Post-Nonlinear (PNL) model (Zhang & Hyvärinen, 2009), in which
@@
The LSNM (Immer et al., 2022) extends ANMs by allowing heteroscedastic noise as follows:

### v2

@@ C.2 Structural causal model identifiability from observational data
where $B\in\mathbb{R}^{d\times d}$ collects the coefficients expressing each $X_{i}$ as a linear function of its parents plus a disturbance $S_{i}$. With mutually independent, non-Gaussian noise terms, the model is identifiable; this is known as the Linear Non-Gaussian Acyclic Model (LiNGAM) (shimizu2006alinear).
@@
An Additive Noise Model (ANM) (hoyer2008anm; peters2014causal) defines each causal variable as a function of (potentially) nonlinear mechanisms and an additive noise contribution:
@@
The most general class with known sufficient conditions for identifiability of the graph is the Post-Nonlinear (PNL) model (zhang2009pnl), in which
@@
The LSNM (immer2022on) extends ANMs by allowing heteroscedastic noise as follows: